Can someone help me on this problem. I am still confused on it.
I need to solve for x and show proof:
-4 (2x-3) = 8x + 5
Thanks,
Sher

Answer by THANApHD(104)

-4 (2x-3) = 8x + 5
-4*2x -(-4)*3 = 8x + 5
-8x +12 = 8x+ 5
-8x +12 - 12 = 8x -8x +5 - 12
-16x = -7
x = -7/-16
= 7/16